OpenStudy (anonymous):

For integral (-5 to -20) for 1/x dx, why is the answer ln(1/4) ?

OpenStudy (anonymous):

What's the derivative of lnx?

OpenStudy (anonymous):

I understand that part, but I don't get how you can just change -5 to -20 to 5 to 20

OpenStudy (anonymous):

ln doesn't exist on the negative side right?

OpenStudy (anonymous):

\[\int\limits_{-5}^{-20} \frac{ 1 }{ x }dx = \int\limits_{5}^{20} \frac{ 1 }{ x }dx\] is this what you're trying to say

OpenStudy (anonymous):

Note that \(\displaystyle \int \frac{1}{x}\,dx = \ln \color{red}|x\color{red}|+C\).

OpenStudy (anonymous):

You flip integrals to make them negative -> positive and vise versa

OpenStudy (anonymous):

the limits of integration

OpenStudy (anonymous):

So it's the absolute signs?

OpenStudy (anonymous):

Yes, they're important to have here.

OpenStudy (anonymous):

\[\int\limits_{a}^{b} f(x) dx = F(b)-F(a) \implies -\int\limits_{a}^{b} f(x) = - (F(b)-F(a))\] but in your case it's absolute value and this is just the fundamental theorem of calculus

OpenStudy (anonymous):

Do you know why we need the absolute values?

OpenStudy (anonymous):

Not really.

ganeshie8 (ganeshie8):

the function 1/x is defined in (-5, -20) so there will be area under/above this curve too. you can find it using integrals or approximating

OpenStudy (anonymous):

oh right, I am finding the area under 1/x, not ln|x|

OpenStudy (anonymous):

You can't evaluate the logarithm without absolute values, I guess if you remember the piece wise function it would be easier to visualize

ganeshie8 (ganeshie8):

you're finding area under/above 1/x and it equals ln|x| as long as you stay away from 0

OpenStudy (anonymous):

Yeah I got confused

OpenStudy (anonymous):

When \(x>0\), \(\ln x\) is defined and thus \(\dfrac{d}{dx}\ln x = \dfrac{1}{x}.\) This then implies the indefinite integral formula \(\displaystyle \int\frac{1}{x}\,dx =\ln x+C\). However, when \(x<0\), \(\ln x\) is no longer defined. We then note that if \(x<0\), then \(-x>0\). Hence, \(\ln (-x)\) is defined for \(x<0\) and we have \(\dfrac{d}{dx}\ln(-x)=\dfrac{1}{-x}\cdot (-1) = \dfrac{1}{x}\). This then implies the indefinite integral formula \(\displaystyle \int \frac{1}{x}\,dx = \ln(-x)+C\). Putting the two pieces together, we see that \[\int\frac{1}{x}\,dx = \begin{cases}\ln x + C & x>0\\ \ln(-x)+C & x<0\end{cases}\] Recalling that \(|x|=\begin{cases}x & x\geq 0\\ -x & x< 0\end{cases}\), we can now see that \(\ln|x|=\begin{cases}\ln x & x>0\\ \ln(-x) & x< 0\end{cases}\). Therefore, \(\displaystyle\int\frac{1}{x}\,dx = \begin{cases}\ln x + C & x>0\\ \ln(-x)+C & x<0\end{cases}=\ln|x|+C\).